hi I can't remeber exactly how to do this problem.
Solve using substitution method:
15x-10y=1
5y=-1+10x
I tried this:
5/y=-1+10/5
then (5)15x-(5)10(5)(-1+10x/5)=1(5)
75x-50(-1+10x)=5
75x+50-500x=5
425x=45
it just dosen't look right

Answer by jim_thompson5910(35256)

5y=-1+10x

2*5y=2*(-1+10x)

10y = -2+20x
-------------------------------------------------------

15x-10y=1

15x-(-2+20x)=1

15x+2-20x = 1

-5x+2 = 1

-5x = 1-2

-5x = -1

x = -1/(-5)

x = 1/5

Now use this to find y.